There are 6 sets of twins among 250 students. What percent of the students have a twin?
Answer by stanbon(75887)

There are 6 sets of twins among 250 students. What percent of the students have a twin?
---
x/100 = 12/250
x = 100(12/250)
x = 24/5
x = 4.2%
===============
Cheers,
Stan H.